OpenStudy (anonymous):

Simplify completely quantity 5 x squared minus 21 x minus 20 all over quantity 5 x squared minus 16 x minus 16

OpenStudy (aroub):

5x^2-21x-20/5x^2-16x-16 ?

OpenStudy (anonymous):

yes :)

OpenStudy (aroub):

Okay :) Well, all you have to do is factor and cancel.

OpenStudy (aroub):

When you factor 5x^2-21x-20 it will become (x-5)(5x+4)

OpenStudy (aroub):

and when you factor 5x^2-16x-16 it will become (5x+4)(x-4)

OpenStudy (anonymous):

so itll be x-5/x-4?

OpenStudy (aroub):

Yes!

OpenStudy (anonymous):

can you help withSimplify completely quantity 2 x squared plus 13 x plus 20 all over x squared minus 5 x minus 36 @aroub

OpenStudy (aroub):

Yeah :)

OpenStudy (aroub):

2x^2+13x+20/x^2-5x-36 Same steps as before, lets factor! 2x^2+13x+20 = (2x+5)(x+4) x^2-5x-36 = (x-9)(x+4)

OpenStudy (aroub):

=> (2x+5)(x+4)/(x-9)(x+4) By cancelling, = 2x+5/x-9

OpenStudy (anonymous):

2x+5/x-9?

OpenStudy (aroub):

=)

OpenStudy (anonymous):

thank you!

OpenStudy (aroub):

No probleemmm!